# Day-Ahead Electricity Price Forecasting Using Merit-Order Curves Time Series

### v1

@@ 6 Results & Discussion
6.1 Functional Principal Component Analysis
The first five and four supply and demand FPCs can be better interpreted on Fig. 6, which displays the effect a perturbation caused by FPCs on the mean curve for a snapshot of the dynamic FPCA at the middle of the test period. In particular for demand, FPC1 explains the quasi-totality of the variability and concerns the magnitude of the total demanded quantity. The contribution to the total variance beyond FPC1 is negligible. Subsequent FPCs were however selected because the elbow method was in this case less conservative than the threshold method. The additional consideration of their temporal consistency highlighted in 5 suggests their significance and goes against the hypothesis that these FPCs just capture noise. It is likely instead that they could capture the small scale variability in a domain of the curve where the intersection with the supply curve is likely to occur, and that is therefore of interest. For supply, FPC1 measures the overall supply level at any price, while FPC2 contrasts situations where there is higher supply at low price (below 100€/MWh) compared to high price (above 200€/MWh) and vice versa. FPC3 and FPC4 regard the distribution of the prices above (resp. below) 180€/MWh, and similarly contrasts situations when these prices are higher or lower. FPC5 is more difficult to interpret and regards more complex features of the price/volume distribution.
@@ 6.2 Curves forecasting
An example of curves (and clearing price) forecast for each of the nine curve-based models can be found in A (Fig. 13). The average squared correlation and the squared correlation function of the nine curves forecasting models are presented in Table 2 and Fig. 7, respectively. The naive model is clearly outperformed by all the other models for both sides. For demand, all models (except naive) perform almost equally: This is very likely due to the low demand elasticity on the Italian day-ahead market meaning that the quantity demand curve is more or less constant equal to the load forecast – predictor to which all models have access to. Concerning supply, the FPCA representation (FPCA) clearly outperforms that of ziel_electricity_2016 (ZST), while the selected variant of the multivariate forecasting model does not show much impact, especially for FPCA. Looking at the entire squared correlation curve, we notice the performance drop of ZST models in the high prices range, which partly explains their lower average .
6.3 Clearing price forecast
@@
The global clearing price forecasting performance is reported on Table 3 with the associated DM tests in Fig. 8(a), for the four price-based and eight curve-based models. All models perform well compared to the naive approach. The DM tests reveal that all ZST models are significantly outperformed by all FPCA, regardless of the multivariate model variant considered. Besides, none of ZST models significantly outperform any of the price-based models, while all of them but one are significantly outperformed by at least one of the price-based models. The FPCA curve-based models instead are the best performers as all of them significantly beat all ZST and price-based approaches151515At the 5% level except for the univariate FPCA models vs fARX for which it is 10%.. A key result is that the consideration of the cross-scores dependence (VARX models) appear to have a significant (positive) impact while the cross-hour dependence (full models) does not show any improvement, whether it be for ZST or FPCA. Instead, for the price-based models, the cross-hour dependence is necessary to reach top performance.
@@ 6.4 Price probabilistic forecasting
Fig. 11 shows the hour-level average CRPS of the best models of each category. As for point forecasting, the best curve-based model performs better in the early afternoon between 12 and 16 than the price-based approach. The performances are nearly equal for the rest of the day.
Looking at the calibration and resolution components (Fig. 14 and 15 left in B), we notice the poor reliability of the curve-based ZST and price-based fARX-IDR models. The first is strongly over-dispersed while the tails of the second are too light. fARX-N forecasts are reliable but show a poor resolution. The other models show comparable calibration – with a slight under-dispersion and clear skew towards lower prices – but heterogeneous resolutions: low for fARX-CP, intermediate for fARX-QRM and higher for FPCA models.
6.5 Computational efficiency
@@
7 Conclusion
Our curve-based model was rigorously compared with that of ziel_electricity_2016 – which share the same multivariate forecasting framework and differs only in the curves representation strategy – and state-of-the-art price-based models for clearing price forecasting. The results show that considering a functional principal component representation brings consequent improvements over a discrete representation and that our curve-based model – even in its simplest form – significantly enhances point and probabilistic clearing price forecasting performance, especially around mid-day when price occasionally drops due to high renewable generation and low demand. Specifically, our curve-based model reduces the mean absolute error by 5% and the continuous ranked probability score by 4%. If we restrict the comparison to the 10:00-17:00 hour range, the reduction is of 10% and 7%, respectively.
Additional analyses could examine the effect of alternative curves representations161616Such as the original form and/or unconstrained forms (c.f. section 2.3)., the vector space dimension, the use of a single versus hour-specific representation and the calibration window size on both curves and clearing price forecasting performance. Besides, more complex non-linear models such as feedforward neural networks could be considered to jointly predict the 24-hour score matrix. Finally, the recent switch to 15-min intervals in European day-ahead markets, making no longer 24 but 96 curves pairs to predict, could justify a functional rather than multivariate perspective to capture intraday dependence. These aspects will be explored in future research.
Acknowledgments
The authors would like to thank Riccardo Barilli and Giulio Caprara of Energy Team S.p.A. for their support in the creation of the dataset used in this analysis and their contribution to the interesting and deep discussions on the topics developed in this paper. This work was funded by Energy Team S.p.A. and MUR, grant Dipartimento di Eccellenza 2023-2027.
Code and data availability
An open-source software implementation of the methods described in this paper, along with the code and data needed to reproduce all numerical results, are available online at https://github.com/gkchln/moc_forecast.
CRedit author statement
Guillaume Koechlin: Conceptualization, Methodology, Software, Writing - Original Draft, Visualization. Filippo Bovera: Conceptualization, Methodology, Writing - Review & Editing. Piercesare Secchi: Conceptualization, Methodology, Writing - Review & Editing.
Appendix A Example of curves and clearing price prediction
Appendix B Reliability and resolution of probabilistic forecasts

### v2

@@ 3 Results & Discussion
We analyze the day-ahead electricity market (MGP) of the Italian Power Exchange (IPEX) over the 2023-2024 period. The testing period covers the whole year 2024 (366 days). Note that probabilistic price predictions require at least 182 days of out-of-sample point predictions so the testing period for probabilistic forecasting starts on July 1st, 2024, covering a period of 6 months. The quantity curves are built according to the procedure described in section 2.1, using the DomandaOfferta dataset retrieved from the FTP server of Gestore del Mercato Elettrico (GME) – the operator of IPEX – which contains all supply offers and demand bids submitted on the MGP. Additionally, we consider a total of 4 exogenous variables for both curves and price forecasts which are similar to those of Mestre et al. (2020) and commonly chosen in EPF: load forecast, wind and solar generation forecast, and forecasts of transfer capacity with neighboring zones (France and Switzerland). Additional details on curves and exogenous data are provided in B.
The analysis of FPCA results and an empirical assessment of computational efficiency are left in C and D, respectively, while an example of curves and clearing price forecast for each of the nine curve-based models is pictured in E (Fig. 13).
@@ 3.1 Curves forecasting
The average squared correlation and the squared correlation function of the nine curves forecasting models are presented in Table 2 and Fig. 3, respectively. The naive model is clearly outperformed by all the other models for both sides. For demand, all models (except naive) perform almost equally: This is very likely due to the low demand elasticity on the Italian day-ahead market meaning that the quantity demand curve is more or less constant equal to the load forecast – predictor to which all models have access to. Concerning supply, the FPCA representation (FPCA) clearly outperforms that of Ziel and Steinert (2016) (ZST), while the selected variant of the multivariate forecasting model does not show much impact, especially for FPCA. Looking at the entire squared correlation curve, we notice the performance drop of ZST models in the high prices range, which partly explains their lower average .
3.2 Price point forecasting
@@
The global clearing price forecasting performance is reported on Table 3 with the associated DM tests in Fig. 4(a), for the four price-based and eight curve-based models. All models perform well compared to the naive approach. The DM tests reveal that all ZST models are significantly outperformed by all FPCA, regardless of the multivariate model variant considered. Besides, none of ZST models significantly outperform any of the price-based models, while all of them but one are significantly outperformed by at least one of the price-based models. The FPCA curve-based models instead are the best performers as all of them significantly beat all ZST and price-based approaches101010At the 5% level except for the univariate FPCA models vs fARX for which it is 10%.. A key result is that the consideration of the cross-scores dependence (VARX models) appear to have a significant (positive) impact while the cross-hour dependence (full models) does not show any improvement, whether it be for ZST or FPCA. Instead, for the price-based models, the cross-hour dependence is necessary to reach top performance.
@@ 3.3 Price probabilistic forecasting
Fig. 7 shows the hour-level average CRPS of the best models of each category. As for point forecasting, the best curve-based model performs better in the early afternoon between 12 and 16 than the price-based approach. The performances are nearly equal for the rest of the day. Reliability and sharpness assessment can be found in F.
4 Conclusion
Our curve-based model was rigorously compared with that of Ziel and Steinert (2016) – which share the same multivariate forecasting framework and differs only in the curves representation strategy – and state-of-the-art price-based models for clearing price forecasting. The results show that considering a functional principal component representation brings consequent improvements over a discrete representation and that our curve-based model – even in its simplest form – significantly improves point and probabilistic clearing price forecasting performance, especially around mid-day when price occasionally drops due to high renewable generation and low demand. Specifically, our curve-based model reduces the mean absolute error by 5% and the continuous ranked probability score by 4%. If we restrict the comparison to the 10:00-17:00 hour range, the reduction is of 10% and 7%, respectively.
Additional analyses could examine the effect of alternative curves representations, the vector space dimension, the use of a single versus hour-specific representation and the calibration window size, on both curves and clearing price forecasting performance. Besides, more complex non-linear models such as feedforward neural networks could be considered to jointly predict the 24 vector representations within a single model. Finally, the recent switch to 15-min intervals in European day-ahead markets, making no longer 24 but 96 curves pairs to predict, could justify a functional rather than multivariate perspective to capture intraday dependence. These aspects will be explored in future research.
Appendix A Multivariate predictive distribution estimation
We estimate the multivariate distribution of the -dimensional forecast error vector . To robustly estimate this distribution which we assume stationary within the calibration window, we consider hour-dependent variances for the marginals but a shared correlation structure: We assume the errors for a given hour are independent and identically distributed (i.i.d.) according to some -dimensional distribution with mean and a covariance matrix which can be decomposed as follows:
Appendix B Details on curves and exogenous data
Relaxation of transmission constraints
Around 35% of the time, the country-level economic optimum is not feasible due to transmission capacity constraints, which means that the market cannot be cleared from the country-level supply and demand curves. In this case, GME runs an optimization procedure, known as locational marginal pricing (LMP), which splits the entire country-level market pool in a minimum number of ”subpools” of market zones such that, in each subpool, the economic optimum is feasible. As a result, there can be several pairs of supply and demand curves – hence several market clearing prices – for a single hour, one for each subpool. To keep the analysis simple, we did as if the country-level economic optimum was always feasible and built the country-level supply and demand curves for any hour, ignoring the problem of transmission congestion. As a result, the constructed curves reflect the true market clearing 65% of the time, the remaining 35% corresponding to a “virtual” market clearing only. Note the same curves construction procedure was considered in Mestre et al. (2020). The associated market clearing price is what GME calls the national price without constraints, and though not always being the exact market price for the different Italian bidding zones, is still a key monitored market indicator.
Market coupling
Restriction of the curves domain
Three example pairs of supply and demand curves of the dataset are showed on Fig. 9.
Exogenous predictors
We consider a total of 4 exogenous variables for both curves and price forecasts which are similar to those of Mestre et al. (2020) and commonly chosen in EPF:
1.
Day-ahead load forecast (1 variable): Country-level day-ahead load forecast obtained from the ENTSO-E Transparency Platform (ENTSO-E, 2025).
2.
3.
Day-ahead forecast of wind and solar generation (1 variable): Combined country-level day-ahead forecasts of wind and solar generation obtained from a private data provider (LSEG Data & Analytics, 2025), derived from the European Centre for Medium-Range Weather Forecasts (ECMWF) and Global Forecast System (GFS) weather models.
Appendix C FPCA results
The first five and four supply and demand FPCs can be better interpreted on Fig. 12, which displays the effect a perturbation caused by FPCs on the mean curve for a snapshot of the dynamic FPCA at the middle of the test period. In particular for demand, FPC1 explains the quasi-totality of the variability and concerns the magnitude of the total demanded quantity. The contribution to the total variance beyond FPC1 is negligible. Subsequent FPCs were however selected because the elbow method was in this case less conservative than the threshold method. The additional consideration of their temporal consistency highlighted in 11 suggests their significance and goes against the hypothesis that these FPCs just capture noise. It is likely instead that they could capture the small scale variability in a domain of the curve where the intersection with the supply curve is likely to occur, and that is therefore of interest. For supply, FPC1 measures the overall supply level at any price, while FPC2 contrasts situations where there is higher supply at low price (below 100€/MWh) compared to high price (above 200€/MWh) and vice versa. FPC3 and FPC4 regard the distribution of the prices above (resp. below) 180€/MWh, and similarly contrasts situations when these prices are higher or lower. FPC5 is more difficult to interpret and regards more complex features of the price/volume distribution.
Appendix D Computational efficiency
@@
Appendix E Example of curves and clearing price prediction
Appendix F Reliability and resolution of probabilistic forecasts
Methodology
Probabilistic forecasting performance is composed of a reliability (or calibration) and sharpness (or resolution111111The two concepts are equivalent when probabilistic forecasts have perfect reliability.) component (Nowotarski and Weron, 2018). As any proper scoring rule, the CRPS presents the advantage of simultaneously evaluating both but it can be interesting to understand how models perform on each of these two components. We assess therefore reliability with the probability integral transform (Gneiting and Katzfuss, 2014)
Results
Looking at the calibration and resolution components (Fig. 14 and 15), we notice the poor reliability of the curve-based ZST and price-based fARX-IDR models. The first is strongly over-dispersed while the tails of the second are too light. fARX-N forecasts are reliable but show a poor resolution. The other models show comparable calibration – with a slight under-dispersion and clear skew towards lower prices – but heterogeneous resolutions: low for fARX-CP, intermediate for fARX-QRM and higher for FPCA models.